OpenStudy (anonymous):

csc(cos^-1 7/25) how do i find this without a calculator?

OpenStudy (bittuaryan):

\[cosec(\cos^{-1} \frac{ 7 }{ 25 } )\] Assume: \[\ \cos^{-1} 7/25 = \theta \] \[\rightarrow \theta=74\] Now, \[cosec\theta =1.040582726326743\]